Adv. Math. Econ. 8, 297-314 (2006)

Variational problem associated with a model of welfare economics with a measure spaces of agents

Alexander Ioffe

1 Department of mathematics,Technion,Haifa 32000,Israel

Received: February 12, 2005 Revised: September 16, 2005 JEL classification: C62 Mathematical Subject Classification (2000): 49J05, 49K05, 49N15, 91B60